OpenStudy (anonymous):

what is the slope intercept form of -2x+5y=15

OpenStudy (aripotta):

can you tell me what slope intercept form is?

OpenStudy (anonymous):

y=mx+b

OpenStudy (triciaal):

this equation has 15 = rearrange the equation to get y = whatever you do to one side of the equation also do to the other to keep the balance

OpenStudy (aripotta):

so we need to get y by itself on the left side. do you know what your first step would be

OpenStudy (anonymous):

yeah, i know that but im really horrible at math and i just dont understand ow to work the problem out.

OpenStudy (aripotta):

-2x + 5y = 15 since -2x is negative, you have to add 2x to both sides and that cancels it out on the left 5y = 2x + 15

OpenStudy (aripotta):

5y = 2x + 15 then divide both sides by 5 to get y by itself y = 2/5x + 3

OpenStudy (aripotta):

does that make sense

OpenStudy (anonymous):

yes thanks so much!!

OpenStudy (aripotta):

no problem :) can you give me a medal please? :}

OpenStudy (anonymous):

how do i do that? im sorry i just joined this site.

OpenStudy (aripotta):

just click on the "best response" button

OpenStudy (anonymous):

oh ok, and thanks for giving me one:)

OpenStudy (aripotta):

no problemoo

OpenStudy (anonymous):

sorry to ask but could you help me with a few more please?

OpenStudy (aripotta):

only if you try them out yourself first!

OpenStudy (anonymous):

i have since my 1st block at 10am this morrning and im clueless lol

OpenStudy (aripotta):

how about you try the next one and tell me what you think the answer is. then i'll tell you if you got it right or not

OpenStudy (anonymous):

ok sounds fair

OpenStudy (anonymous):

ok the problem is -x - 6y +30 and i got y = -6/1 +31

OpenStudy (aripotta):

there's no = in your problem tho. and there's no x in your solution. i think you made some typos

OpenStudy (anonymous):

ya sorry, its -x -6y = 30

OpenStudy (aripotta):

-x -6y = 30 so you add x to both sides -6y = x + 30 then divide both sides by -6 y = -1/6x - 5

OpenStudy (anonymous):

oh alright, well i was close

OpenStudy (aripotta):

if you need more help, open a new thread and then tag me on the question

OpenStudy (anonymous):

ok thank u